License plates in a particular state have two letters, three single-digit numbers, and then one letter. If no plate can begin with the letter ‘A’ and all three of the letters must be different, how many different license plates are possible?
Answer by edjones(8007)

LL-NNN-L
.
1st two letters: 25*26=650
.
3 numbers: 10^3=1,000
.
Last letter: 26
.
650*1000*26=16,900,000
.
Ed